Let f(z Find the following functions. Simplify your answers. f(g(x)) = g(f(x)) = an r-5 Preview Preview
Answer #1

x f (x)-- , g(x) =

1. f(g(x)) = \frac{g(x)}{g(x)-5}

5 5

  = \frac{\frac{5}{x}}{\frac{5-5x}{x}}

z (5- 5)

  (5 - 5.x)

So f(g(x)) = \frac{5}{(5-5x)}

2. g(f(x)) = \frac{5}{f(x)}

= \frac{5}{\frac{x}{x-5}}

= \frac{5*(x-5)}{x}

5r -25